Shift report 20010731 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jul 31 2001 - 00:15:39 EDT

  • Next message: Apache: "Shift report 20010731 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jul 31 00:15:38 2001, Shift: 16:00-24:00
    Supervisor: KH, with: R. Karabowicz
    Log book#6 pages 120 - 124
    
    Shift Summary:
    ------------------------------------------------------------
    Beam delivered at 23:45
    Stored and cogged
    Started run 3941
    
    Achievements:
    P. C. diagnosed the problem with H1 TDC.  In the end he
    swapped out the TDC and the problem went away.  This was
    diagnosed by generating a simple ROOT tree and looking at
    correlations.  He learned that the problem channels were
    directly correlated with each other indicating cross talk.
    
    Problems:
    No new problems that have not been discussed before.
    
    Machine status:
    Spent nearly the whole shift on machine development.  Some
    problems with the RF during a part of the shift.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jul 31 2001 - 00:16:27 EDT